OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5 a)-4x - 9y = -19 -10y = -30 b)4x + 3y = 5 2y = -6 c)7x - 3y = -11 9x = -6 D)12x - 3y = -33 14x = -28

OpenStudy (anonymous):

a) 4x - y = -11 2x + 3y = 5 -y=-11-4x /:(-1) y=11+4x 2x+3(11+4x)=5 2x+33+12x=5 14x+33=5/-33 14x=-28/:14 x=-2 4*2-y=-11 9-y=-11/-9 -y=2/:(-1) y=-2